If f is continuous on [a,b]t hen \underline_{\int_a^b}f(x)dx=\overline{\int_a^b}f(x)dx
1
Expert's answer
2020-11-01T16:43:26-0500

Solution


Since f(x) is continuous on [a,b] then it is uniformly continuous on the interval.

Let "\epsilon>0" by uniform continuity there exist a "\delta>0" such that "|f(x)-f(y)|< \epsilon" whenever "|x-y|<\delta" such that "x,y\in [a,b]"

We now divide the interval [a,b] into partitions "J_1,J_2,\dots,J_n" each of lengths "\frac{b-a}{N}"

By definition

"\underline{\int_a^bf(x)}dx\leq \sum_{k=1}^N\left(\inf_{x \in J_k}f(x)\right)|J_k|"


"\overline{\int_a^bf(x)}dx\leq \sum_{k=1}^N\left(\sup_{x \in J_k}f(x)\right)|J_k|"

"\overline{\int_a^bf(x)}dx-\underline{\int_a^bf(x)}dx\leq \sum_{k=1}^N\left(\sup_{x \in J_k}f(x)-\inf_{x \in J_k}f(x)\right)|J_k|" ...... (1)

However we have

"|f(x)-f(y)|<\epsilon" "\forall x,y \in J_k" ,since "|J_k|=\frac{b-a}{N}" ,In particular we have

"f(x)<f(y)+\epsilon" "\forall x,y\in J_k"

Taking suprema in x we obtain

"\sup_{x\in J_k}f(x) \leq f(y) + \epsilon" "\forall y\in J_k"

And taking infima in y we obtain

"\sup_{x\in J_k}f(x) \leq \inf_{y\in J_k} f(y) + \epsilon"

Inserting this into (1) we have

"\overline{\int_a^bf(x)}dx-\underline{\int_a^bf(x)}dx\leq \sum_{k=1}^N\epsilon|J_k|"


"\overline{\int_a^bf(x)}dx-\underline{\int_a^bf(x)}dx\leq(b-a)\epsilon"


But "\epsilon >0" is arbitrary and (b-a) is fixed. Thus "\overline{\int_a^bf(x)}dx-\underline{\int_a^bf(x)}dx" cannot be positive, but "\overline{\int_a^bf(x)}dx\geq\underline{\int_a^bf(x)}dx" which implies "\overline{\int_a^bf(x)}dx-\underline{\int_a^bf(x)}dx=0" which complete the proof.